OpenStudy (anonymous):

Will give medal and be a fan. Given f(x)=x^2+3 and g(x)=x+5/x. find (gof)(4). !

OpenStudy (anonymous):

okay,so do you want it graphed?

OpenStudy (anonymous):

Yeah , the answer i got is 14/19 but i think im wrong

OpenStudy (anonymous):

okay, I'm doing this stuff so I think I can figure it out:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know what x eaquals?

OpenStudy (anonymous):

equals*

OpenStudy (anonymous):

Okay, how did you get 14/19? can you show me your work?

OpenStudy (anonymous):

wait is it 15.. or -15.. iajsodijaso. No.. :/ ive been working on math for the past 2 hours so..sorry brain farts.. >.<

OpenStudy (anonymous):

Yeah, I have to.. hate it. Homeschooled?

OpenStudy (anonymous):

I'm sorry, I am having a hard time figuring it out as well. Sorry:/

OpenStudy (anonymous):

Yeappp

OpenStudy (anonymous):

._. so am i right or no

OpenStudy (anonymous):

ya there?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

(gof)(4) is short for first find f(4) can you find f(4) ?

OpenStudy (phi):

f(x)=x^2+3 f(4) is short for, "replace x with 4, and simplify"

OpenStudy (anonymous):

oh okay..i kinda get it...

OpenStudy (phi):

what do you get for f(4) ?

OpenStudy (anonymous):

Sorry im writing it down. is it 15? <

OpenStudy (anonymous):

@phi

OpenStudy (phi):

f(x)=x^2+3 are you saying when x is replace with 4, like this f(4)= 4^2 +3 you get 15 ? is your calculator broken?

OpenStudy (phi):

(gof)(4) is short for first find f(4) can you find f(4) ? if you can, make that number x in g(x) g(x)=x+5/x to find (g o f)(4)

OpenStudy (anonymous):

no but ive been doing this work for like 2 1/2 hours.. this is the last one

OpenStudy (phi):

what do you get for f(4) ?

OpenStudy (phi):

what do you get for f(4) ? f(4)= 4^2 +3 what is 4*4 ? and then add 3. what do you get ?